# Mousavi Wang Ravanelli Subakan

### v1

@@ Abstract
Large Language Models (LLMs) are widely used in Spoken Language Understanding (SLU). Recent SLU models process audio directly by adapting speech input into LLMs for better multimodal learning. A key consideration for these models is the cross-modal alignment between text and audio modalities, which is a telltale sign as to whether or not LLM is able to associate semantic meaning to audio segments. While various methods exist for fusing these modalities, there is no standard metric to evaluate alignment quality in LLMs. In this work, we propose a new metric, ALAS (Automatic Latent Alignment Score). Our study examines the correlation between audio and text representations across transformer layers, for two different tasks (Spoken Question Answering and Emotion Recognition). We showcase that our metric behaves as expected across different layers and different tasks.
keywords:
Large Language Models, Multimodal LLMs, Cross-modal alignment
@@ 1 Introduction
Large Language Models (LLMs) [1, 2, 3, 4, 5] have recently demonstrated very strong performance in text-based tasks, often reaching human-level capabilities in natural language processing (NLP). A key factor contributing to this success is instruction tuning [6, 7, 8, 9, 10], where models are trained on instruction-response pairs to improve their ability to follow open-ended prompts.
There has been increasing interest in multimodal learning, where LLMs are extended to process diverse inputs such as images, speech, and audio [11, 12, 13, 14, 15, 16, 17, 18, 19]. These models typically integrate a pretrained modality-specific encoder with a connection (adapter) module that fuses different input representations. The alignment process is crucial because it ensures that different modalities are mapped into a shared representation space, allowing the model to extract meaningful cross-modal relationships. Training such models often involves cross-modal pretraining and instruction tuning to improve their ability to process multimodal data effectively.
Alignment plays a fundamental role in the success of these multimodal models, particularly in speech-language understanding (SLU). Tan et al. [20] explored various alignment strategies for optimizing the connection module and showed that stronger alignment leads to improved model performance. Effective alignment ensures that spoken and textual representations convey the same semantic information, reducing ambiguity and improving model performance. Several approaches have been developed to address the speech-text alignment problem. The Montreal Forced Aligner (MFA) [21], built on the Kaldi framework [22], provides phoneme-level alignment using forced alignment techniques. Connectionist Temporal Classification (CTC) [23] is widely used for speech-text alignment, as it enables alignment without explicit labels [24, 25]. More recently, models such as UnitY2 [26] and ZMM-TTS [27] have introduced unit-based alignment techniques using pretrained encoders, demonstrating improved performance across multiple languages.
Another important strategy for improving alignment in multimodal LLMs involves training with large-scale aligned speech-to-text datasets. Many models adopt a multi-stage training approach [13, 11, 28, 29], where the first stage is dedicated to learning high-quality alignment between auditory and textual representations, mostly using large-scale automatic speech recognition or audio captioning datasets. This is followed by fine-tuning for specific downstream tasks. However, there is no standardized metric to evaluate how well an SLU model learns speech-text alignment. Measuring alignment quality is critical for understanding whether an LLM effectively captures semantic relationships between speech and text. A well-defined alignment metric would provide deeper insights into cross-modal learning and help guide improvements in model design.
In this work, our contributions are as follows,
  • We propose ALAS (Automatic Latent Alignment Score), a new metric for assessing alignment quality in speech-text multimodal LLMs.
@@
  • We apply ALAS on two different tasks (Spoken Question Answering and Emotion Recognition) and analyze the level of alignment between the text and audio modalities.
@@ 2.1 Automatic Pipeline for Latent Alignment Score
In order to measure the alignment score, we need an aligned test dataset, meaning that for each audio sample, we must have the corresponding transcription. We begin by selecting an appropriate dataset and defining the assigned task. We use speech question answering (LibriSQA dataset [30]), and emotion recognition (IEMOCAP dataset [31]) in this paper.
@@
Step 4: Measuring the Alignment Score. Given three inputs—text latent representations, audio latent representations, and the reference alignment from Whisper—we compute the final alignment score using our scoring module. Our scoring module consists of two sub-modules. As we discuss next in more detail, we first measure the cross-modal cosine similarities, and then apply monotonic-alignment-search algorithm [32] on the cross-modal cosine similarity matrices.
@@ 2.3 Automatic Latent Alignment Score (ALAS)
After obtaining the similarity score matrix, $S^{l}\in[0,1]^{T\times A}$, we obtain the alignment path between the two modalities using the monotonic-alignment-search algorithm [32], where $T$ and $A$ respectively denote the total number of time steps respectively in the text and audio modalities.
In order then to assess the quality of the alignments that we estimate through similarities and monotonic-alignment-search, we obtain word-level alignments from Whisper [^1] https://huggingface.co/openai/whisper-large-v3 [33]. The final score is defined such that,
@@ 3 Experiments
We evaluate ALAS with two popular multimodal LLMs: SALMONN-7B[11] [^2] https://huggingface.co/tsinghua-ee/SALMONN-7B/tree/main and Qwen2-Audio-Instruct-7B[13] [^3] https://huggingface.co/Qwen/Qwen2-Audio-7B-Instruct (We refer to these models as SALMONN and QWEN for brevity). These models have achieved state-of-the-art (SOTA) results on various audio question-answering tasks. Both models leverage multitask learning and multistage training to improve alignment between modalities.
@@
We experiment with two speech datasets that contain paired audio and text. The first dataset is LibriSQA[30], which is designed for free-form and open-ended Speech Question Answering (SQA). Given an audio clip or its corresponding transcription, we prompt the model with a question in text format and evaluate its response. The second dataset is IEMOCAP[31], which is used for emotion recognition. The task involves classifying the input into one of four emotion categories: happy, sad, angry, and neutral. The choice of these datasets allows us to analyze the alignment patterns of multimodal speech LLMs in two different types of tasks. LibriSQA requires the model to extract semantic meaning from speech, making it a useful benchmark for evaluating how well the model aligns spoken and textual representations in tasks that rely on linguistic understanding. In contrast, emotion recognition focuses more on acoustic features rather than linguistic content, providing insight into how these models process non-semantic auditory cues. By evaluating the alignment score on both datasets, we can examine how well the alignment mechanism generalizes across tasks that pertain to different aspects of speech. In Figure 4 in the Appendix, we illustrate the prompts used for both QWEN and SALMONN models.
For the Spoken Question-Answering Task, to measure response similarity, we use Sentence-BERT[34] [^4] https://huggingface.co/sentence-transformers, where we pass each generated response through the sentence transformer to compute a similarity score. We discard the pairs with a similarity score lower than 0.7 to ensure that both text and audio contain comparable task-relevant information. We observe that QWEN and SALMONN respectively obtains scores of 0.71 and 0.76 respectively on the SQA task. For the Emotion Recognition task, the models respectively obtain scores of 21.56%, and 49.30% [35] on the IEMOCAP dataset.
@@ 3.1 Qualitative Analysis
In Figure 2, we qualitatively compare the cross-modal similarity patterns obtained with Qwen2-Audio-Instruct-7B (left) and Salmonn-7B (right), for the LibriSQA dataset (SQA task). We showcase the similarity $S^{l}$ for layers 0, 5, 10, 32, where layer 0 represents the output after applying a linear transformation, just before the transformer layers of the LLMs. We observe that at layer 0, the similarity patterns exhibit much more vague diagonal-maximal-paths compared to the crisp diagonal-similarity patterns obtained for later layers, which suggests that alignment between modalities become more distinguishable on the later transformer layers. That is the say, we observe a much clear diagonal-like path for later layers (e.g layer 32), compared to the earlier layers (e.g. layer 0, layer 5). Note that layer 0 denotes the representations right after the feature extractor, and layer 1 is the output of the first transformer layer of the LLM.
In Figure 2, we also showcase the monotonic alignment search algorithm applied on the similarity patterns. We show the Whisper alignment path with a white trace (estimated by using the word boundaries estimated by Whisper), and we use a red trace for the alignment path found after applying the monotic search algorithm on the similarity matrix $S^{l}$ for layer $l$ (this is the path that is used to compute ALAS, which we call the ALAS-path). We observe that for QWEN, earlier layers indicate more significant deviation from the ground truth path (approximated by Whisper). Especially on layer 5, we observe that the deviation is very visible. However with later layers, we observe that the ALAS-paths get distinctly more similar the Whisper-path. We quantitatively verify these observations in the next section.
@@ 3.2 Quantitative Analysis
In Figure 3, we show the average-ALAS score with respect to different layers, for QWEN and SALMONN on LibriSQA and IEMOCAP. We observe that for the LibriSQA task, the later layers generally tend to exhibit better ALAS score, as noted for the qualitative comparisons as well, in Section 3.1.
We note that at layer 0 (before the transformer layers) the ALAS score is a bit better compared to the immediately subsequent transformer layers of the LLM. This is not very surprising since Qwen2-Audio-Instruct-7B, and SALMONN-7B both uses Whisper-v3 [33] as the feature encoder, which outputs semantically informed representations. The application of subsequent transformer layers initially projects these representations that decreases the similarity as shown by the plots, however, the semantic alignment quickly catches up with more layers, and later even decreases further, for instance for QWEN on the LibriSQA dataset. For SALMONN on the LibriSQA dataset, we observe that the alignments closely follow the reference Whisper alignments in absolute sense. This is reasonable since the temporal resolution of SALMONN on the audio domain is lower compared to QWEN.
Another interesting observation that arises from this quantitative analysis is that, unlike the question-answering task on the LibriSQA dataset, for the Emotion Recognition task the level of alignment actually decreases for the last few layers. This is in fact reasonable since Emotion Recognition is not necessarily a semantic task. For certain recordings the speakers express their emotions through speech patterns and prosody, and not necessarily through the semantic content of the audio. This observation therefore suggests that ALAS is able to capture the reduction in semantic alignment between the audio and text modalities for different tasks.
@@ 4 Conclusions
Semantic alignment between different modalities is an important property for a multimodal LLM, which indicates that the LLM is indeed able to learn semantic relationships between text and the audio modality. Measuring the semantic alignment is therefore an important problem to assess multi-modal LLMs. In this paper, we propose a new metric (ALAS) for measuring the semantic alignment between the audio and text modalities of multimodal speech LLMs. We have tested this metric for two different tasks within the domain of spoken language understanding (SLU). We have analyzed the ALAS scores qualitatively and quantitatively for two popular multimodal LLMs, for different layers and we observe that ALAS is able to pick-up on expected patterns. These observed patterns include, increasing alignment with respect to layers on a more semantic task like question-answering, and decreasing alignment on a less semantic task such as emotion recognition. We will release our code upon acceptance of this paper to allow researchers to easily evaluate their multi-modal LLMs.

### v2

# ALAS: Measuring Latent Speech-Text Alignment For Spoken Language Understanding In Multimodal LLMs

@@ Abstract
Large Language Models (LLMs) are increasingly used in Spoken Language Understanding (SLU), where effective multimodal learning depends on the alignment between audio and text. Despite various fusion methods, no standard metric exists to assess this alignment. This work introduces ALAS (Automatic Latent Alignment Score), a metric that evaluates alignment by measuring correlations between audio and text representations across transformer layers. Experiments on Spoken Question Answering and Emotion Recognition show that ALAS captures meaningful patterns across tasks and layers.
Audio LM, Speech Understanding, Alignment
@@ 1 Introduction
Large Language Models (LLMs) (Touvron et al., 2023; Chowdhery et al., 2023; Brown et al., 2020; Du et al., 2021; Achiam et al., 2023) have shown impressive performance on text-based NLP tasks, often achieving human-level results. A major contributor to this success is instruction tuning (Peng et al., 2023; Zhang et al., 2023b; Taori et al., ; Li et al., 2023; Zhang et al., 2023a), which improves a model’s ability to follow open-ended prompts. Building on these advances, recent efforts have extended LLMs to multimodal inputs like images, speech, and audio (Tang et al., 2023; Chen et al., 2023; Chu et al., 2024; Wu et al., 2023; Li et al., 2023; Dai et al., 2023; Rubenstein et al., 2023; Zhan et al., 2024; Xie & Wu, 2024; Mousavi et al., 2025b). These models typically pair a pretrained modality-specific encoder with an adapter module to fuse inputs, and rely on alignment to ensure representations from different modalities are projected into a shared semantic space. Training often combines cross-modal pretraining with instruction tuning to enhance multimodal understanding.
Alignment is particularly crucial in speech-language understanding (SLU). Tan et al. (Tan et al., 2024) demonstrated that optimizing alignment in the adapter improves performance, as it enables audio and text to convey consistent semantic information. Traditional alignment methods include Montreal Forced Aligner (MFA) (McAuliffe et al., 2017), based on Kaldi (Povey et al., 2011), and CTC-based approaches (Graves et al., 2006; Sainath et al., 2020; Huang et al., 2024), which allow label-free alignment. More recent unit-based methods, such as UnitY2 (Barrault et al., 2023) and ZMM-TTS (Gong et al., 2024), use pretrained encoders to improve multilingual alignment performance.
Many multimodal LLMs adopt a multi-stage training setup (Chu et al., 2024; Tang et al., 2023; Hu et al., 2024; Gong et al., 2023; Mousavi et al., 2025a), where an initial stage focuses on learning robust audio-text alignment from large ASR or audio captioning datasets, followed by task-specific fine-tuning. Despite the importance of alignment, no standard metric exists to evaluate how well a model learns these cross-modal correspondences. Such a metric is critical for understanding whether an LLM truly captures semantic relationships across modalities. This work addresses that gap with the following contributions:
  • We propose ALAS (Automatic Latent Alignment Score), a new metric for assessing alignment quality in speech-text multimodal LLMs. Our method is model-and task-agnostic, requires no manual transcription or forced alignment, and can be applied to any dataset.
@@ 2.1 Automatic Pipeline for Latent Alignment Score
In order to measure the alignment score, we need an aligned test dataset, meaning that for each audio sample, we must have the corresponding transcription. We begin by selecting an appropriate dataset and defining the assigned task. We use speech question answering (LibriSQA dataset (Zhao et al., 2023)), and emotion recognition (IEMOCAP dataset (Busso et al., 2008)) in this paper.
@@
Step 4: Measuring the Alignment Score. Given three inputs—text latent representations, audio latent representations, and the reference alignment from Whisper—we compute the final alignment score using our scoring module. Our scoring module consists of two sub-modules. As we discuss next in more detail, we first measure the cross-modal cosine similarities, and then apply monotonic-alignment-search algorithm (Kim et al., 2020) on the cross-modal cosine similarity matrices.
@@ 2.3 Automatic Latent Alignment Score (ALAS)
After obtaining the similarity score matrix, $S^{l}\in[0,1]^{T\times A}$, we obtain the alignment path between the two modalities using the monotonic-alignment-search algorithm (Kim et al., 2020), where $T$ and $A$ respectively denote the total number of time steps respectively in the text and audio modalities. In order then to assess the quality of the alignments that we estimate through similarities and monotonic-alignment-search, we obtain word-level alignments from Whisper(Radford et al., 2022). The final score is defined such that,
@@ 3 Experiments
We evaluate ALAS with two popular multimodal LLMs: SALMONN-7B(Tang et al., 2023) and Qwen2-Audio-Instruct-7B(Chu et al., 2024) (We refer to these models as SALMONN and QWEN for brevity). These models have achieved state-of-the-art (SOTA) results on various audio question-answering tasks. Both models leverage multitask learning and multistage training to improve alignment between modalities. We experiment with two speech datasets that contain paired audio and text. The first dataset is LibriSQA(Zhao et al., 2023), which is designed for free-form and open-ended Speech Question Answering (SQA). Given an audio clip or its corresponding transcription, we prompt the model with a question in text format and evaluate its response. The second dataset is IEMOCAP(Busso et al., 2008), which is used for emotion recognition. The task involves classifying the input into one of four emotion categories: happy, sad, angry, and neutral. The choice of these datasets allows us to analyze the alignment patterns of multimodal speech LLMs in two different types of tasks. LibriSQA requires the model to extract semantic meaning from speech, making it a useful benchmark for evaluating how well the model aligns spoken and textual representations in tasks that rely on linguistic understanding. In contrast, emotion recognition focuses more on acoustic features rather than linguistic content, providing insight into how these models process non-semantic auditory cues. By evaluating the alignment score on both datasets, we can examine how well the alignment mechanism generalizes across tasks that pertain to different aspects of speech. In Figure 4 in the Appendix, we illustrate the prompts used for both QWEN and SALMONN models.
@@
For the Spoken Question-Answering Task, to measure response similarity, we use Sentence-BERT(Reimers & Gurevych, 2019), where we pass each generated response through the sentence transformer to compute a similarity score. We discard the pairs with a similarity score lower than 0.7 to ensure that both text and audio contain comparable task-relevant information. We observe that QWEN and SALMONN respectively obtain scores of 0.71 and 0.76,respectively on the SQA task. For the Emotion Recognition task, the models respectively obtain scores of 21.56%, and 49.30% (Wang et al., 2024) on the IEMOCAP dataset.
@@ 3.1 Qualitative Analysis
In Figure 2, we qualitatively compare the cross-modal similarity patterns obtained with Qwen2-Audio-Instruct-7B (left) and Salmonn-7B (right), for the LibriSQA dataset (SQA task). We showcase the similarity $S^{l}$ for layers 0, 5, 10, 32, where layer 0 represents the output after applying a linear transformation, just before the transformer layers of the LLMs. We observe that at layer 0, the similarity patterns exhibit much more vague diagonal-maximal-paths compared to the crisp diagonal-similarity patterns obtained for later layers, which suggests that alignment between modalities become more distinguishable on the later transformer layers. That is the say, we observe a much clear diagonal-like path for later layers (e.g layer 32), compared to the earlier layers (e.g. layer 0, layer 5). Note that layer 0 denotes the representations right after the feature extractor, and layer 1 is the output of the first transformer layer of the LLM. In Figure 2, we also showcase the monotonic alignment search algorithm applied on the similarity patterns. We show the Whisper alignment path with a white trace (estimated by using the word boundaries estimated by Whisper), and we use a red trace for the alignment path found after applying the monotic search algorithm on the similarity matrix $S^{l}$ for layer $l$ (this is the path that is used to compute ALAS, which we call the ALAS-path). We observe that for QWEN, earlier layers indicate more significant deviation from the ground truth path (approximated by Whisper). Especially on layer 5, we observe that the deviation is very visible. However with later layers, we observe that the ALAS-paths get distinctly more similar the Whisper-path. We quantitatively verify these observations in the next section.
@@ 3.2 Quantitative Analysis
In Figure 3, we show the average-ALAS score with respect to different layers, for QWEN and SALMONN on LibriSQA and IEMOCAP. We observe that for the LibriSQA task, the later layers generally tend to exhibit better ALAS score, as noted for the qualitative comparisons as well, in Section 3.1. We note that at layer 0 (before the transformer layers) the ALAS score is a bit better compared to the immediately subsequent transformer layers of the LLM. This is not very surprising since Qwen2-Audio-Instruct-7B, and SALMONN-7B both uses Whisper-v3 (Radford et al., 2022) as the feature encoder, which outputs semantically informed representations. The application of subsequent transformer layers initially projects these representations that decreases the similarity as shown by the plots, however, the semantic alignment quickly catches up with more layers, and later even decreases further, for instance for QWEN on the LibriSQA dataset. For SALMONN on the LibriSQA dataset, we observe that the alignments closely follow the reference Whisper alignments in absolute sense. This is reasonable since the temporal resolution of SALMONN on the audio domain is lower compared to QWEN. Another interesting observation that arises from this quantitative analysis is that, unlike the question-answering task on the LibriSQA dataset, for the Emotion Recognition task the level of alignment actually decreases for the last few layers. This is in fact reasonable since Emotion Recognition is not necessarily a semantic task. For certain recordings the speakers express their emotions through speech patterns and prosody, and not necessarily through the semantic content of the audio. This observation therefore suggests that ALAS is able to capture the reduction in semantic alignment between the audio and text modalities for different tasks.
@@ 4 Conclusions
Semantic alignment between different modalities is an important property for a multimodal LLM, which indicates that the LLM is able to learn semantic relationships between text and the audio modality. Measuring the semantic alignment is therefore an important problem to assess multi-modal LLMs. We propose a new metric (ALAS) for measuring the semantic alignment between the audio and text modalities of multimodal speech LLMs. We have analyzed the ALAS scores qualitatively and quantitatively for two popular multimodal LLMs, for different layers. ALAS captures meaningful patterns across transformer layers: alignment increases in semantic tasks like question answering and decreases in less semantic tasks like emotion recognition.